Be sure to answer all parts.

Calculate the Ka of a weak acid if a 0.063 M solution of the acid has a pH of 2.85 at 25

°

C.

Ka =

×

10
(Enter your answer in scientific notation.)
Answer #1

+ HA C c (1-2) Ž H+ 0 ca To G Á 0 ca a pH=2.85 [ut] = 10-2.85 = 1.41% 10-3 M Lut = ca =0.063M xa = 1.41 x los Q = 0.0224 [HA]